OpenStudy (anonymous):

1. A game is said to pay “3 to 1” if for every $1 bet you win, you get paid $3. In roulette, there are 40 spaces on the wheel: 19 red, 19 black, and 2 green. Red and black pay 2 to 1 and green pays 20 to 1. if I bet 10$ on red for 6 games, what can I expect to win (or lose)? I know I need to use the expected value formula just not sure how to set it up

jimthompson5910 (jim_thompson5910):

not exactly sure how roulette is played, but do you have to land on red every time (all 6 times) to win if you bet on red?

OpenStudy (anonymous):

would it be (-10)(1)+(2)(19/40)+(3)(19/40)+(4)(19/40)+(5)(19/40)+(6)(19/40)=

OpenStudy (anonymous):

no I don't think so

OpenStudy (anonymous):

that's wrong because it say for every 1$ bet you win 3$

jimthompson5910 (jim_thompson5910):

well if you're just focusing on one game, and you bet $10 on red, the expected winnings are E[X] = P(winning)*V(Winning) + P(losing)*V(losing) E[X] = (19/40)*(20) + (21/40)*(-10) E[X] = 9.5 + (-5.25) E[X] = 4.25

jimthompson5910 (jim_thompson5910):

not 100% sure, but it seems like it's as simple as saying 6*(E[X]) = 6*(4.25) = 25.5 would be your expected winnings after 6 games...but then again, that might be too high

OpenStudy (anonymous):

I'll let you know

jimthompson5910 (jim_thompson5910):

alright thanks

OpenStudy (anonymous):

(-$10)(21/40) + ($30)(19/40) -5.25 + 14.25 =$9.00 $9.00 *6 = $54.00

jimthompson5910 (jim_thompson5910):

I thought red paid 2 to 1 (not 3 to 1)?

OpenStudy (anonymous):

it say 3 to 1 odds

jimthompson5910 (jim_thompson5910):

hmm at the very top it says "Red and black pay 2 to 1"

OpenStudy (anonymous):

oh wait I see what you are talking about

jimthompson5910 (jim_thompson5910):

there might be a typo? idk

OpenStudy (anonymous):

(-10)(21/40) + (20)(19/40) -5.25 + 9.50 = $3.75 *6 = 22.50 that sounds better?

jimthompson5910 (jim_thompson5910):

(-10)(21/40) + (20)(19/40) = 4.25 though

OpenStudy (anonymous):

yep you are right thank you I would of gotten it wrong twice

jimthompson5910 (jim_thompson5910):

you're welcome